OpenStudy (anonymous):

Between which two consecutive whole number does the square root of 38 lie

OpenStudy (accessdenied):

Would you agree that \( \sqrt{42} \) is greater than \(\sqrt{38} \)? And \(\sqrt{26}\) is less than \(\sqrt{38} \)? We are looking for two perfect squares underneath the radical so that \(\sqrt{a^2} = a \) is a whole number.

OpenStudy (accessdenied):

So rather than think about how to find the whole numbers, let's try to find the two perfect squares nearest to 38. There is one greater than it and one smaller than it. The square root will lie between the squared value of each. (e.g. 64 is the perfect square, 8 is the squared value i refer to)

OpenStudy (anonymous):

6 ,7. 6squared is 36 and 7 squared is 42 therefore 38 lies between these two numbers

OpenStudy (ipwnbunnies):

^^ There is your answer young turtle. Good job.

OpenStudy (ipwnbunnies):

\[\sqrt(36) < \sqrt(38) < \sqrt(49)\] \[6 < \sqrt(38) < 7\]